In a distribution where X=21 and S=3, what is the z-score?
Answer and rationalization:
1
Expert's answer
2020-06-10T19:56:37-0400

"Z = \frac {X-M} {S}"


"Z = \frac {21-M} {3}"

Assume M = 20

then Z score

"Z = \frac {21-20} {3}=0.33"

Z(0.33) = 0.62930